OpenStudy (anonymous):

rewrite 6tan3x in terms of tan x

OpenStudy (kc_kennylau):

Welcome to OpenStudy :D You may want to read the code of conduct ( http://openstudy.com/code-of-conduct ). Don't forget to click on the button called "Best Answer" to give a medal to the person who helped you :)

OpenStudy (praxer):

@Nicole15 did you get your result.

OpenStudy (anonymous):

no @praxer

OpenStudy (praxer):

tan(a+b) = [tan(a) + tan(b)]/[1 - tan (a) * tan (b)] tan(2x+x) = [tan2x + tanx]/[1 - tan2x * tanx] ........ (1) tan2x = tan(x+x) = [tanx + tanx]/[1 - tanx * tanx] = 2tanx/(1-(tanx)^2) ........ (2) Making (2) in (1): tan(3x) = [2tanx/(1-(tanx)^2) + tanx]/[1 - 2tanx/(1-(tanx)^2) * tanx] tan(3x) = [(2tanx + tanx - (tanx)^3)/((1-(tanx)^2))]/[1 - (2(tanx)^2)/(1-(tanx)^2)] tan(3x) = [(3tanx - (tanx)^3)/(1-(tanx)^2)]/[(1-(tanx)^2 - 2(tanx)^2)/(1-(tanx)^2)] tan(3x) = [(3tanx - (tanx)^3)/(1-(tanx)^2)]/[(1-3(tanx)^2)/(... \[tan(3x) = [3tanx - (tanx)^3]/[1-3(tanx)^2]\] now you have the tan3x formula. So for 6tan3x just multiply a 6 with the result. Understood it @Nicole15 :)